OpenStudy (anonymous):

I really need help! I need someone to explaint it too! 3(x + 1) = −2(x − 1) − 4 solve for x

OpenStudy (anonymous):

I tried so hard and still cant get it. help?

OpenStudy (anonymous):

ok you just need to simplify the equation by steps, Start on the right side: and take 3(x+1), multiply the 3 through, so you get 3x+3. then the other side, take -2(x-1) - 4 and multiply that -2 through. Now you have -2x +2 -4. Simplify that again and you get -2x - 2. Now rewrite the whole thing. 3x+3=-2x-2. Put the x's on the left so subtract the -2x across, therefore 5x+3=-2, then take the +3 to the right half, so 5x= -5. now divide each side by 5, and you get X= -1. Hope that helps.

OpenStudy (anonymous):

thx so much the first three steps i got it but got confused with the others

OpenStudy (anonymous):

one more question plz

OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4